A motorboat can go 10 miles downstream in 20 minutes. It takes 30 minutes for this boat to go back upstream the same 10 miles. Find the speed of the current.
Not sure if I did this one right, I got 10 mph. Is this right?
Thanks!!

Answer by Paul(988)

20(x+10)=30(x-10)
20x+200=30x-300
20x=30x-500
-10x=-500
x=50
Hence, the speed of the current is 50mph
Check: 20[(50)+10]=30[(50)-10]
1200=1200
L.S.=R.S.